OpenStudy (anonymous):

Solve each proportion: 4.) x/4=13/52 5.) x/2x+1=16/40 6.) 9/10=9x/70 7.) 2/7=b+1/56 8.) 11/y=9/27 9.) 3/34=m/51 PLEASE HELP I HAVE NO IDEA WHAT TO DO!!!! ):

OpenStudy (campbell_st):

well these take a little cross multiplying in 4. x = (13*4)/52 multiply by the denominator of the left hand fraction. 5. 40x = 16(2x + 1)... or 40x = 32x + 16 cross multiply then solve for x 6. multiply by the denominator on the right (9*70)/10 = 9x then solve for x hope this gives you the idea.

OpenStudy (anonymous):

you have to put them in proportion

OpenStudy (anonymous):

hold on lemme do them

OpenStudy (anonymous):

Okay!(:

OpenStudy (anonymous):

4. 1

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

4. x=1

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

all you really had to do is 4*13=x*52

OpenStudy (anonymous):

5, 6, and 7 are super confusing to me

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

For 8. I got 33 and for 9. I got 4.5

OpenStudy (anonymous):

6=x=7

OpenStudy (anonymous):

I need to show work.

OpenStudy (anonymous):

6. 9*70=630 divided by 10 =63 now you divide 63/9=7 do you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats so funny @Luis_Rivera and @beautiful_ladie how did you do 8?

OpenStudy (anonymous):

The answer was given correctly in the first solution, every single one, you just cross multiply. 1) 52*x=13*4, therefore x=1 and the ration is 1/4 2) 40*x=(2x+1)*16, therefore x=2, the ratio is 2/5 3) 9*70=9*x*10, therefore x=7, the ratio is 63/70 4) 2*56=(b+1)*7, therefore b=15, and the ratio is 16/56 5) 11*27=9*y, therefore y=33, and the ratio is 11/33 6) 3*51=34*m, therefore m=4.5 and the ratio is 4.5/51

OpenStudy (anonymous):

crossed multiplied and got 9y=297, then divided by 9

OpenStudy (anonymous):

correct @Khushal_Shah

OpenStudy (anonymous):

@beautiful_ladie you got 8 and 9 right

OpenStudy (anonymous):

How about 7?

OpenStudy (anonymous):

it is 15 look at what @Khushal_Shah wrote

OpenStudy (anonymous):

I only see up to 6

OpenStudy (anonymous):

a good website you can go on for these problem is www.mathway.com

OpenStudy (anonymous):

hey, i started at 1 instead of 4. just add three to all my numbers and you'll get the right answers